OpenStudy (anonymous):

2tan x / 1-tan^2 x = 1/√3. What does x equal? a. 11pi/12 + n(pi) b. 5pi/12 + n(pi) c. 7pi/12 + n(pi) d. pi/12 + n(pi)

OpenStudy (anonymous):

\[\tan 2x=\frac{ 2 \tan x }{ 1-\tan ^2 x }\]

OpenStudy (anonymous):

wait what..?

OpenStudy (anonymous):

can you proceed further?

OpenStudy (anonymous):

No, I don't get any of this. I am so confused.

OpenStudy (anonymous):

\[\tan 2x=\frac{ 1 }{ \sqrt{3} }=\tan \frac{ \pi }{ 6 }=\tan \left( n \pi+\frac{ \pi }{ 6 } \right)\],where n is an integer.

OpenStudy (anonymous):

okay..